amaya123:

If a, b, c, and d are four different numbers and the proportion 2004-02-04-01-00_files/i0290000.jpg is true, which of the following is false?

Arieonna:

any answer choices

amaya123:

a,5/a = d/c b, a/c = b/d c, b/a = c/d d, a+b/c= c+d/d

Arieonna:

@krissy525 can help u i hope

amaya123:

thanks

Krissy525:

@arieonna wrote:
@krissy525 can help u i hope
My brain hurts

Krissy525:

@astrid1

DeadOfWinter:

Nothing here to support the question. Unanswerable.

DeadOfWinter:

I see it say, JPG right? Is that an image?

jhonyy9:

@amaya123 please try share this image 2004-02-04-01-00_files/i0290000.jpg use the Attach File button below to this window